(x-1)/3 = (x+1)/5
How do I solve this step by step?

Answer by checkley77(12569)

(x-1)/3 = (x+1)/5
5(X-1)=3(X+1)
5X-5=3X+3
5X-3X=3+5
2X=8
X=8/2
X=4 ANS.
PROOF:
(4-1)/3=(4+1)/5
3/3=5/5
1=1